# Language Arithmetics: Towards Systematic Language Neuron Identification and Manipulation

### v2

@@ 5.1 Identification Insights
The layer-wise distribution of language-specific neurons in Llama-3.1, Mistral Nemo, and Aya-Expanse partially aligns with prior findings Tang et al. (2024); Kojima et al. (2024); Zhao et al. (2024). We observe that these neurons are predominantly concentrated in the later layers, whereas Tang et al. (2024) and Kojima et al. (2024) report a dual-peaked distribution in older models, with concentrations in both early and late layers. Notably, Tang et al. (2024) also show that the more multilingual BLOOM model Le Scao et al. (2023) exhibits a similar late-layer concentration, a trend further supported by Mondal et al. (2025) with recent multilingual models. This divergence likely reflects architectural differences, or variations in training data composition. Interestingly, Chou et al. (2025) demonstrate that modifying a small number of SAE features from mid-to-late layers enables high-accuracy language steering for high-resource languages, while Andrylie et al. (2025) similarly find that their SAE-LAPE method identifies language-specific features concentrated in the same layers. These converging results, obtained through entirely different methodologies, reinforce our observation that language-specific representations are predominantly localized in the mid-to-late transformer layers of modern multilingual LLMs.
Layer-wise logit lens analysis complements these findings: in all tested models, English token probabilities emerge early on and persist, while target language tokens appear more sharply in the very final layers, which is in line with Wendler et al. (2024) and Schut et al. (2025). Entropy trends in more English-centric models rise toward the end, suggesting that language generation decisions are finalized in the later layers. In terms of cross-linguistic neuron overlap, we observe clear clustering patterns based on language families. This supports findings from Tan et al. (2024), who report that related languages share substantial neural components while maintaining distinct processing features.
@@ 5.2 Manipulation Insights
Our findings show that neuron-level interventions can reliably steer multilingual model behavior, with our additive language arithmetic method outperforming activation replacement Tang et al. (2024); Kojima et al. (2024) and DiffMean-based approaches Marks and Tegmark (2023). Activating language-specific neurons improves performance across both generative and discriminative tasks, with the strongest gains observed when the activated language matches the target or is typologically related, yielding positive transfer within language families, but more limited benefits for structurally distant languages. Interestingly, Hindi and Urdu show disproportionate improvements across many interventions, potentially due to neurons encoding multiple overlapping functions (Elhage et al., 2022). Moreover, compared to Mondal et al. (2025), who report limited improvements from activation replacement or language neuron fine-tuning on tasks like XNLI Conneau et al. (2018a) and XQuAD Artetxe et al. (2019), our method achieves consistent gains due to three key differences. First, we evaluate across a broader and more typologically diverse set of languages. Second, our additive interventions using mean activation vectors preserve more of the original representational context. Third, all evaluations are conducted in a strict zero-shot setting, without fine-tuning on downstream tasks in any source language.
The effects are especially pronounced in our language forcing experiments. Deactivating neurons for the input language while simultaneously activating neurons for a desired target language significantly increases the likelihood of generating output in the target language–even when the input and output languages are typologically distant. These results suggest that target-language neurons contain sufficient signal to override source-language priors, especially for languages better represented in pre-training. In contrast, low-resource languages are harder to control, likely due to weaker or less distinct neural representations resulting from limited pretraining exposure. Importantly, manipulation also reveals internal "fallback" mechanisms: when dominant language neurons like English or French are deactivated, the model switches to secondary high-resource languages such as Russian or Arabic. This "fallback" hierarchy varies slightly across models but consistently exposes the model’s internal language preferences. The persistence of certain languages despite deactivation–particularly English–highlights the robustness of high-priority language priors in LLMs.
@@ Acknowledgments
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Federal Ministry of Research, Technology and Space (BMFTR) as part of the project TRAILS (01IW24005).
@@ Appendix B Selected Languages
Language Code Language Name
bo Tibetan
mt Maltese
it Italian
es Spanish
de German
ja Japanese
ar Arabic
zh Chinese
af Afrikaans
nl Dutch
fr French
pt Portuguese
ru Russian
ko Korean
hi Hindi
tr Turkish
pl Polish
sv Swedish
da Danish
no Norwegian
en English
Table 2: Language codes and corresponding full names
@@ Appendix C Neuron Counts
Appendix D Downstream Baseline Results
Model af ar bo da de es fr hi it ja ko mt nl no pl pt ru sv tr zh
Llama-3.1 0.227 0.049 0.172 0.233 0.176 0.154 0.276 0.113 0.155 0.024 0.040 0.086 0.136 0.163 0.100 0.311 0.166 0.217 0.096 0.102
Mistral-Nemo 0.114 0.022 0.176 0.109 0.140 0.144 0.215 0.023 0.123 0.140 0.022 0.030 0.088 0.076 0.039 0.190 0.104 0.069 0.041 0.136
(a) Machine translation (BLEU scores).
Model af ar bo da de en es fr hi it ja ko mt nl no pl pt ru sv tr zh
Llama-3.1 0.690 0.760 0.316 0.742 0.764 0.847 0.751 0.764 0.670 0.766 0.729 0.767 0.542 0.750 0.744 0.714 0.777 0.799 0.787 0.727 0.779
Mistral-Nemo 0.689 0.751 0.284 0.706 0.792 0.841 0.823 0.790 0.357 0.806 0.749 0.798 0.408 0.733 0.697 0.659 0.783 0.840 0.714 0.734 0.808
(b) Machine comprehension (accuracy).
Model ar bg de el en es fr hi ru sw th tr ur vi zh
Llama-3.1 0.382 0.361 0.365 0.356 0.429 0.358 0.377 0.289 0.375 0.350 0.386 0.397 0.130 0.358 0.379
Mistral-Nemo 0.332 0.333 0.335 0.335 0.338 0.333 0.349 0.331 0.335 0.332 0.334 0.335 0.327 0.334 0.334
(c) Natural language inference (accuracy).
Model ar de el en es hi ro ru th tr vi zh
Llama-3.1 0.410 0.462 0.422 0.616 0.515 0.460 0.542 0.416 0.459 0.511 0.606 0.355
Mistral-Nemo 0.281 0.443 0.258 0.585 0.556 0.387 0.451 0.362 0.379 0.452 0.444 0.276
(d) Question answering (F1 score).
Table 7: Baseline multilingual evaluation results for translation, comprehension, NLI, and QA.
Appendix E Language Neuron Distributions
@@
Figure 10: Distribution of individual language-specific neurons across Llama-3.1 layers for all 21 languages.
@@ Appendix F Additional Results for Aya Models
Appendix G Output Analysis with Logit Lens
@@
Appendix H Language Forcing
@@
(b) Activate-only - Top 2%
(c) Activate-only - Top 3%
(d) Deactivate-Activate - Top 1%
(e) Deactivate-Activate - Top 2%
(f) Deactivate-Activate - Top 3%
Figure 26: Page 1 - Language forcing results across 5 neuron ratios for two manipulation strategies - activate-only and deactivate-activate for Llama-3.1-8B. For deactivation, we set the neurons to 0.
(a) Activate-only - Top 4%
@@
Figure 27: Page 2 - Language forcing results across 5 neuron ratios for two manipulation strategies - activate-only and deactivate-activate for Llama-3.1-8B. For deactivation, we set the neurons to 0.
Appendix I Language "Fallbacks"

### v3

@@ 5.1 Identification Insights
The layer-wise distribution of language-specific neurons in Llama-3.1, Mistral-Nemo, and Aya-Expanse partially aligns with prior findings Tang et al. (2024); Kojima et al. (2024); Zhao et al. (2024). We observe that these neurons are predominantly concentrated in the later layers, whereas Tang et al. (2024) and Kojima et al. (2024) report a dual-peaked distribution in older models, with concentrations in both early and late layers. Notably, Tang et al. (2024) also show that the more multilingual BLOOM model Le Scao et al. (2023) exhibits a similar late-layer concentration, a trend further supported by Mondal et al. (2025) with recent multilingual models. This divergence likely reflects architectural differences or variations in pre-training data composition and training procedures. Interestingly, Chou et al. (2025) demonstrate that modifying a small number of SAE features from mid-to-late layers enables high-accuracy language steering for high-resource languages, while Andrylie et al. (2025) similarly find that their SAE-LAPE method identifies language-specific features concentrated in the same layers. These converging results, obtained through entirely different methodologies, reinforce our observation that language-specific representations are predominantly localized in the mid-to-late transformer layers of modern multilingual LLMs.
Layer-wise logit lens analysis complements these findings: in all tested models, English token probabilities emerge early on and persist, while target language tokens appear more sharply in the very final layers, which is in line with Wendler et al. (2024), Schut et al. (2025), and Wang et al. (2025). Entropy trends in more English-centric models rise toward the end, suggesting that language generation decisions are finalized in the later layers. In terms of cross-linguistic neuron overlap, we observe clear clustering patterns based on language families. This supports findings from Tan et al. (2024), who report that related languages share substantial neural components while maintaining distinct processing features.
@@ 5.2 Manipulation Insights
Our findings show that neuron-level interventions can reliably steer multilingual model behavior, with our additive language arithmetic method outperforming activation replacement Tang et al. (2024); Kojima et al. (2024) and DiffMean-based approaches Marks and Tegmark (2023). Activating language-specific neurons improves performance across both generative and discriminative tasks, with the strongest gains observed when the activated language matches the target or is typologically related, yielding positive transfer within language families for some tasks, but more limited benefits for structurally distant languages. Interestingly, Hindi and Urdu show disproportionate improvements across many interventions, potentially due to neurons encoding multiple overlapping functions (Elhage et al., 2022). Moreover, compared to Mondal et al. (2025), who report limited improvements from activation replacement or language neuron fine-tuning on tasks like XNLI Conneau et al. (2018a) and XQuAD Artetxe et al. (2019), our method achieves performance gains due to three key differences. First, we evaluate across a broader and more typologically diverse set of languages. Second, our additive interventions using mean activation values preserve more of the original representational context. Third, all evaluations are conducted in a strict zero-shot setting, without fine-tuning on downstream tasks in any source language.
The effects are especially pronounced in our language forcing experiments. Deactivating neurons for the input language while simultaneously activating neurons for a desired target language significantly increases the likelihood of generating output in the target language–even when the input and output languages are typologically distant. These results suggest that target-language neurons contain sufficient signal to override source-language priors, especially for languages better represented in pre-training. In contrast, low-resource languages are harder to control, likely due to weaker or less distinct neural representations resulting from limited pretraining exposure. Importantly, manipulation also reveals internal "fallback" mechanisms: when dominant language neurons like English or French are deactivated, the model switches to secondary high-resource languages such as Russian or Arabic. This "fallback" hierarchy varies slightly across models and language deactivation orders but consistently exposes the model’s internal language preferences. The persistence of certain languages despite deactivation–particularly English–highlights the robustness of high-priority language priors in LLMs.
@@ Acknowledgments
This research was supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a project funded by Horizon Europe under GA No.101079164, by lorAI - Low Resource Artificial Intelligence, a project funded by the European Union under GA No.101136646, and by the German Federal Ministry of Research, Technology and Space (BMFTR) as part of the project TRAILS (01IW24005).
@@ Appendix B Selected Languages
Language Code Language Name Language Family
bo Tibetan Sino-Tibetan
mt Maltese Afro-Asiatic
it Italian Indo-European (Romance)
es Spanish Indo-European (Romance)
de German Indo-European (Germanic)
ja Japanese Japonic
ar Arabic Afro-Asiatic (Semitic)
zh Chinese Sino-Tibetan
af Afrikaans Indo-European (Germanic)
nl Dutch Indo-European (Germanic)
fr French Indo-European (Romance)
pt Portuguese Indo-European (Romance)
ru Russian Indo-European (Slavic)
ko Korean Koreanic
hi Hindi Indo-European (Indo-Aryan)
tr Turkish Turkic
pl Polish Indo-European (Slavic)
sv Swedish Indo-European (Germanic)
da Danish Indo-European (Germanic)
no Norwegian Indo-European (Germanic)
en English Indo-European (Germanic)
Table 2: Language codes, names, and families of the 21 languages used in our experiments. Language family classification follows Glottolog (Hammarström et al., 2024).
@@ Appendix C Neuron Counts
Appendix D Language Neuron Distributions
Figure 10: Distribution of identified language-specific neurons across Llama-3.1-8B layers for all 21 evaluated languages. The neuron distributions for individual languages are further in the Appendix.
@@ Appendix D Language Neuron Distributions
Appendix E Output Analysis with Logit Lens
@@
Appendix F Language Forcing
@@
Figure 27: Detailed language forcing results across 2 neuron ratios for two manipulation strategies - activate-only and deactivate-activate for Llama-3.1-8B. For deactivation, we set the neurons to 0.
Model Intervention Strategy Top 1% Top 2% Top 3% Top 4% Top 5%
Mistral-Nemo Additive Activate 17.49% 18.93% 19.99% 22.45% 22.75%
Additive Deactivate + Activate 25.54% 27.51% 29.59% 32.39% 33.22%
Replacement Activate 12.43% 12.69% 13.86% 15.19% 14.47%
Replacement Deactivate + Activate 20.03% 21.08% 22.22% 24.75% 25.88%
DiffMean Activate 14.51% 16.18% 16.55% 17.15% 18.14%
DiffMean Deactivate + Activate 21.96% 23.69% 25.62% 27.36% 28.11%
Table 7: Overall success rates (%) of language forcing for Mistral-Nemo-12B using three intervention types and two manipulation strategies across different top-k% neuron thresholds.
Model Intervention Strategy Top 1% Top 2% Top 3% Top 4% Top 5%
Aya-Expanse-8B Additive Activate 66.40% 77.48% 81.29% 82.58% 84.73%
Additive Deactivate + Activate 75.59% 83.30% 85.22% 84.47% 86.81%
Replacement Activate 49.62% 61.19% 65.31% 65.61% 69.39%
Replacement Deactivate + Activate 65.72% 73.96% 73.99% 74.11% 77.06%
DiffMean Activate 62.85% 70.52% 75.43% 77.82% 80.88%
DiffMean Deactivate + Activate 73.74% 78.87% 81.67% 82.20% 84.24%
Table 8: Overall success rates (%) of language forcing for Aya-Expanse-8B using three intervention types and two manipulation strategies across different top-k% neuron thresholds.
Appendix G Language "Fallbacks"